OpenStudy (anonymous):

Nick wants to buy a pair of shoes. The original cost of the shoes is $56.75, and the markup is 12 percent. How much will he have to pay for the shoes?

OpenStudy (alexandervonhumboldt2):

he will have to pay 56.75*(1+12/100)=63.56

OpenStudy (anonymous):

Step 1: Multiply the price of the shoes (56.75) by the markup percent (.12) Step 2: Take the number you get from 56.75 x .12 and add it to 56.75 \[56.75 \times .12 = 6.81 \] \[6.81 + 56.75 = 63.56\]